The average time it takes a group of Senior High School students to complete a certain examination is 48.2 minutes. The standard deviation is 7 minutes. If 45 randomly selected college student take the examination, what is the probability that the mean time it takes the group to complete the test between 46 and 49 minutes?


1
Expert's answer
2022-04-14T03:27:04-0400

We have a normal distribution,ย "\u03bc=48.2,\u03c3=7,n=45."

Let's convert it to the standard normal distribution,

"\bar{z}=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}},\\
\bar{z}_1=\cfrac{46-48.2}{7/\sqrt{45}}=-2.11,\\
\bar{z}_2=\cfrac{49-48.2}{7/\sqrt{45}}=0.77,\\
P(46<\bar{X}<49)=P(-2.11<\bar{Z}<0.77)=\\
=P(\bar{Z}<0.77)-P(\bar{Z}<-2.11)=\\
=0.7794-0.0174=0.7620\text{ (from z-table).}"

โ€‹